OpenStudy (anonymous):

On a hot day where the temperature is 32 deg celsius, the distance between two lamp post on the road is 30m as measured by a metal tape whose coefficient of linear expansion is hypothetically (50 x 10^-6)/deg celsius. If the tape gives its correct reading at 20 deg celsius, what is the actual distance between the lamp posts?

OpenStudy (irishboy123):

grabbed this from Wiki \(\dfrac{\Delta L}{L} = \alpha\Delta T\) and you know that \(\Delta L + L = 30 = L(\alpha \Delta T + 1 ) \) solve for L with \(\alpha = 50 \times 10^{-6}\) deg celsius and \(\Delta T = 12\)

OpenStudy (shamim):

We can solve it in little different way

OpenStudy (shamim):

We know the equation L=L0(1+alpha*theta)

OpenStudy (shamim):

If L=30m then Theta=32°c

OpenStudy (shamim):

Then the equation will b 30=L0(1+50*10^-6*32)

OpenStudy (shamim):

We will get another equation for 20°c

OpenStudy (shamim):

L=L0(1+50*10^-6*20)

OpenStudy (shamim):

Then divide equation 2 by equation 1

OpenStudy (shamim):

L=?